# ToxiCraft:A Novel Framework for Synthetic Generation of Harmful Information

### v1

@@ Abstract
In different NLP tasks, detecting harmful content is crucial for online environments, especially with the growing influence of social media. However, previous research has two main issues: 1) a lack of data in low-resource settings, and 2) inconsistent definitions and criteria for judging harmful content, requiring classification models to be robust to spurious features and diverse. We propose Toxicraft, a novel framework for synthesizing datasets of harmful information to address these weaknesses. With only a small amount of seed data, our framework can generate a wide variety of synthetic, yet remarkably realistic, examples of toxic information. Experimentation across various datasets showcases a notable enhancement in detection model robustness and adaptability, surpassing or close to the gold labels. We release the generated data at Github upon acceptance.
Warning: This paper has instances of hateful and offensive language to serve as examples.
@@
Zheng Hui1,2, Zhaoxiao Guo1,3, Hang Zhao1, Juanyong Duan1, Congrui Huang1 1Microsoft Corporation, 2Columbia University, 3Tsinghua University Correspondence: juaduan@microsoft.com
@@ 1 Introduction
The emergence of the Transformer-based modelVaswani et al. (2017) led to the development of complex models capable of identifying toxic content with remarkable accuracy. However, the effectiveness of these models hinges on the quality and diversity of the datasets used for their trainingBanko and Brill (2001). Traditional datasets, often curated manually, tend to lack the breadth required to cover the multifaceted nature of harmful content. Consequently, while these models excel at recognizing explicit instances of toxicity, they struggle when faced with subtler formsCasula and Tonelli (2023).
@@
Numerous datasets containing harmful content are commonly sourced from platforms like Twitter or online forums Davidson et al. (2017); Grimminger and Klinger (2021a); Waseem and Hovy (2016); Sachdeva et al. (2022). Nevertheless, many of these datasets suffer from significant class imbalances regarding specific types of toxic language. Training a compact model tailored for detecting specific harmful information or opting for a smaller model due to resource or latency constraints often necessitates extensive human-labeled data. However, such data are scarce in numerous downstream tasks and costly to annotate Breitfeller et al. (2019); Juuti et al. (2020). These smaller or low-resource datasets also encounter challenges, such as limited linguistic diversity and the risk of overfitting Klubička and Fernández (2018). Privacy concerns arise regarding the use of social media data obtained without user consent in research experiments Casula and Tonelli (2023). Moreover, there is the issue of dataset decay, particularly with online messages, including abusive ones, which tend to be deleted over time Casula and Tonelli (2023) [^1] Twitter’s policy only permits displaying post IDs instead of the text itself, resulting in inaccessible content over time, often due to deletion by either the user or Twitter itself, particularly in settings related to harmful information..
Recent advancements in large language models (LLMs) such as GPT-3Brown et al. (2020) have spurred researchers to delve into their potential for synthetic dataYoo et al. (2021); Ye et al. (2022b); Gao et al. (2023). Although LLM-based data augmentation often excels in enhancing model performance, there are mixed findings regarding whether LLM-generated synthetic data can effectively help model training to achieve performance levels comparable to models trained on real-world data with human annotationLi et al. (2023).
@@
LLMs have also shown their potential in generating synthetic data for harmful information detection. However, the work utilizing LLMs in this area remains limitedSen et al. (2023) and the existing works are relatively outdatedJuuti et al. (2020); Geet d’Sa et al. (2021); Liu et al. (2020); Geet d’Sa et al. (2021); Wullach et al. (2021). Additionally, the effectiveness of LLMs in this area is constrained due to their struggles with subjective tasks and the amplification of biases inherent in human-annotated training data. These biases arise from the varying standards used for ‘gold’ labels in annotation, and LLMs often fail to capture the complex nuances and diversity of human communication in tasks with high subjectivity.
@@
To address these issues, we introduce the ToxiCraft Framework that enhances the robustness and reduces bias in synthetic data by incorporating a diverse set of training examples and employing advanced bias detection algorithms. The ToxiCraft Framework is specifically designed to produce higher quality synthetic data that better reflect human judgements, making models trained on this data more effective and equitable for real-world applications. An example of the flow are showing in Figure 1. In Figure 1, (A) is direct training where a smaller model is trained on a limited dataset, potentially including harmful data, that may affect model quality. (B) is using large language models to segregate harmful from non-harmful data, refining the dataset for training smaller models. (C) is Zero-shot Synthesis which uses GPT-4 to create a synthesized dataset via zero-shot prompts, used to train smaller models without further data. The ‘Ours’ ToxiCraft Method is Utilizing crafted prompts and attributes, in which GPT-4 generates a tailored dataset for training refined smaller models. A more detailed description of our approach can be found in Section 3.
To summarise, our work has the following key contributions:
@@
  • Enhances four low-resource harmful information datasets to address the pervasive issue of harmful dataset sources decay, making the enhanced datasets publicly available to ensure reproducibility and broader utility.
@@
  • Applications of the synthesized data in fine-tuning smaller models, namely BERT and Roberta, highlight the practical value of LLM generated data.
@@ 3 Methodology
In this section, we commence by defining the problem statement and subsequently introduce the ToxiCraft framework. ToxiCraft is designed as a versatile and effective few-shot learning framework that operates through dataset generation. The stages of the ToxiCraft framework are illustrated in Figure 2.
@@ 3.1 Problem Statement
Robustness: Evaluated by the model’s performance consistency across variations in the test sets, specifically those generated by different noise vectors $\epsilon$ within the Toxicraft framework:
$\text{Robustness}(\theta,\epsilon)=\text{Var}_{\epsilon}(\ell(h_{\theta}(x_{\epsilon}),y))$
where $x_{\epsilon}$ are instances generated with specific perturbations, and Var represents variance, measuring performance variability. We also provide a pseudocode in Appendix A.
@@ 3.2 Toxicraft Framework
Toxicraft Prompt Next, we added manually-set indicators to this data with Attribute Extraction by GPT (We call this process as "Toxicraft Prompt") and input 10% of the data extracted randomly from the seed data pool together with "Toxicraft Prompt" to GPT. Manually-set indicators include the intensification or weakening of tone, whether to increase swear words, whether to use irony, country and time. Each value will be randomly selected or masked. Thus, it is possible that none of the manually-set indicators are added, but at most only one answer will be selected from each indicator. For example, the country could be randomly set as the United States and the year as 2023.
@@
Through this carefully designed series of steps, Toxicraft effectively utilizes the Chain of Thought (COT) concept to gradually generate high-quality training data. Additionally, we compared the dataset generated by Toxicraft with the dataset generated only using simple COT prompts, and detailed experimental results will be presented in Section 4.
@@ 4.1 Datasets and processing
In this section, we outline the data employed in our experiments. We first introduce the four datasets containing harmful information in Section 4.1.1. Following this, we consolidate these datasets and present statistics on the unified dataset in Section 4.1.2.
@@ 4.1.1 Harmful information Datasets
Sadly, it’s important to mention many datasets such as AIRAIT Waseem and Hovy (2016), only their tweet IDs and their labels were publicly available. Regrettably, a substantial portion of tweets (>= 40% positive datas) were inaccessible through the X API.
@@ 4.1.2 Data processing and Splits
Our data preprocessing pipeline prioritizes identifying harmful content in English. Thus, we exclude any tweets in languages other than English and verify language using a Roberta-based language identifier. Additionally, to prevent overlap between similar datasets, we remove near-duplicated entries through normalization, ignoring duplicate entries based on their lemmatized forms and removing URLs and mentions. The majority of collected datasets focus on hate speech, and in some cases, offensive speech, which we consider harmful. For datasets with a binary hate classification task or a more detailed classification like CMS, where all "hate" subclasses are treated as one, we categorize them as harmful. Datasets focusing on specific types of hate speech, such as sexism, are also considered harmful. Moreover, datasets containing offensive speech are likewise classified as harmful in our case. Finally, in datasets like AHSD and HASOC, where a distinction between hate, offensive, or profound speech exists, we include entries labeled as hate speech or offensive, considering them as harmful. Entries labeled as normal or not-hateful are categorized as not harmful.
@@
Table 2: Comparison of Model Performances Across Different Databases and Data Generation Methods. The comparison parameter is MacroF1. Results outperformed training on all gold label are underlined while the best performance on each dataset are bolded. Some baseline results come from Antypas and Camacho-Collados (2023).
@@ 4.3 Implementation Details
We employed GPT-4 OpenAI et al. (2024) as our agent model, for generations, in our experiment, we set N to be 1000, K to be 100, and M to be 3. For smaller-scale classification tasks, we utilized two language models of a reasonable size: BERT-base Devlin et al. (2019b) and RoBERTa-base Liu et al. (2019) as downstream models. The implementations provided by Hugging Face Wolf et al. (2020) were utilized for training and evaluating all language models. We fine-tuned the learning rate, warm-up steps, number of epochs, batch size, and other hyperparameters for each model. For BERT, we use learning rate=$1.8282\times 10^{-5}$, epoch=3, batch=4, warm-up step=30.And for RoBERTa, we use learning rate=$1.1856\times 10^{-5}$, epoch=3, batch=4, warm-up step=30. All fine-tuning experiments were conducted using A40 and GTX 4080TI GPU, and the results are reported in terms of macro-F1 score.
@@ 5 Experimental Results
Analysis on Robustness [^2] Due to limited resources, only two different datasets were selected for cross-dataset robustness analysis. Figure 3 reveals Toxicraft’s robust ability to generate synthetic data that enhances model generalization across datasets. In the AHSD to HASOC evaluation, we observe that as the seed size increases, the performance of models trained on Toxicraft synthetic data rises significantly, peaking at a Macro F1 score of 67 for the 200 seed size. This surpasses the gold label performance for the AHSD dataset, which stands at a Macro F1 score of 54. Similarly, for the HASOC to HTPO evaluation, the upward trend continues with the performance of the 200 seed size reaching a Macro F1 score of 58, again outperforming the gold label performance, benchmarked at 56. These results not only demonstrate Toxicraft’s capacity to produce high-quality synthetic data that captures the complexity of harmful content but also its effectiveness in adapting to different datasets. The ability to exceed gold label performance suggests that Toxicraft could potentially reduce the need for extensive labeled datasets, particularly in specialized or sensitive areas like political content, where data acquisition is challenging.
@@ 7 Future Work
Future research for Toxicraft will focus on three main objectives: enhancing multilingual capabilities, refining the seed data selection process, and reducing reliance on high-cost LLMs.
Multilingual Data Generation: Future efforts will explore translating harmful content from English into other languages while preserving both the local nuances and the target language characteristics.
Optimizing Seed Data Selection: Another area of focus will be to refine the methodology for selecting seed data. We aim to develop strategies to identify the most diverse and representative examples within a dataset. One proposed method is to analyze word embeddings and select seed data points whose vectors are the least similar, ensuring a broad coverage of the data space.
Cost-effective Model Exploration: We will also continue our efforts to identify more cost-effective and transparent alternatives to GPT-4, such as the Mistral model, for generating high-quality synthetic samples of harmful content. These steps will potentially lower the barriers to accessing sophisticated data augmentation techniques and contribute to the advancement of in content moderation.

### v2

@@ Abstract
In different NLP tasks, detecting harmful content is crucial for online environments, especially with the growing influence of social media. However, previous research has two main issues: 1) a lack of data in low-resource settings, and 2) inconsistent definitions and criteria for judging harmful content, requiring classification models to be robust to spurious features and diverse. We propose ToxiCraft, a novel framework for synthesizing datasets of harmful information to address these weaknesses. With only a small amount of seed data, our framework can generate a wide variety of synthetic, yet remarkably realistic, examples of toxic information. Experimentation across various datasets showcases a notable enhancement in detection model robustness and adaptability, surpassing or close to the gold labels. Dataset is available at this GitHub repository.
[…]
0.3pt $\blacktriangle$ !OcFFL2exThis paper has instances of hateful and offensive language to serve as examples.
@@
[…]
@@ 1 Introduction
The emergence of the Transformer-based model Vaswani et al. (2017) led to the development of complex models capable of identifying toxic content with remarkable accuracy. However, the effectiveness of these models hinges on the quality and diversity of the datasets used for their training Banko and Brill (2001). Traditional datasets, often curated manually, tend to lack the diversity required to cover the multifaceted nature of harmful content. Consequently, while these models excel at recognizing explicit instances of toxicity, they struggle when faced with subtler forms Casula and Tonelli (2023).
@@
Many datasets containing harmful content are typically sourced from social media platforms like Twitter or online forums Davidson et al. (2017); Grimminger and Klinger (2021a); Waseem and Hovy (2016); Sachdeva et al. (2022). However, these datasets often exhibit significant class imbalances, particularly concerning specific types of toxic language. Training a model to detect specific harmful content or opting for smaller models due to resource or latency constraints requires substantial amounts of human-labeled data. Unfortunately, such data are scarce in downstream tasks and expensive to annotate Breitfeller et al. (2019); Juuti et al. (2020). Additionally, smaller or low-resource datasets face challenges like limited linguistic diversity and a higher risk of overfitting Klubička and Fernández (2018). Privacy concerns also emerge when using social media data obtained without user consent for research purposes Casula and Tonelli (2023). Furthermore, dataset decay is also a notable issue, especially with online social media, including abusive content, which are often deleted over time Casula and Tonelli (2023) [^1] Twitter’s policy only allows the display of post IDs rather than the text itself, making content inaccessible over time, often due to deletion by either the user or Twitter, particularly in cases involving harmful information..
@@
Recent advancements in large language models (LLMs) such as the GPT series Brown et al. (2020) have prompted researchers to explore their potential for generating synthetic data Yoo et al. (2021); Ye et al. (2022b); Gao et al. (2023). While LLM-based data augmentation often improves model performance, findings are mixed regarding whether synthetic data generated by LLMs can consistently match the effectiveness of models trained on real-world, human-annotated data Li et al. (2023). LLMs have also shown promise in generating synthetic data for training harmful content detection models. However, work in this area remains limited Sen et al. (2023) and much of the existing research is relatively outdated Juuti et al. (2020); Geet d’Sa et al. (2021); Liu et al. (2020); Geet d’Sa et al. (2021); Wullach et al. (2021). Moreover, the effectiveness of LLMs in this domain is hindered by their difficulty with subjective tasks and their tendency to amplify biases inherent in human-annotated training data. These biases stem from inconsistent standards used in defining ‘gold’ labels, and LLMs often struggle to capture the complex nuances and diversity of human communication in highly subjective tasks.
@@
To address these challenges, we introduce the ToxiCraft Framework, which enhances robustness and reduces bias in synthetic data generation. The ToxiCraft Framework is specifically designed to produce higher-quality synthetic data that better supports model training. An example of the process is illustrated in Figure 1. In Figure 1, (A) is direct training where a smaller model is trained on a limited dataset, that may affect model quality. (B) shows the use of large language models to do harmful detaction. (C) illustrates Zero-shot Synthesis, where use LLMs to generate a synthetic dataset via zero-shot prompts, used to train smaller models without additional data. Our ToxiCraft method (D) employs carefully crafted prompts and attributes, utilizing GPT-4 as the backbone to generate a synthetic dataset for training more refined smaller models. A more detailed explanation of our approach can be found in Section 3. To summarise, our work has the following key contributions:
@@
  • Enhances low-resource harmful information datasets to address the pervasive issue of harmful dataset sources decay, making the enhanced datasets publicly available to ensure reproducibility and broader utility.
@@
  • Applications of the synthesized data in fine-tuning smaller models, highlight the practical value of LLM generated data.
@@ 3 Methodology
In this section, we commence by defining the problem statement and subsequently introduce the ToxiCraft framework. ToxiCraft is designed as a versatile and effective few-shot learning framework that operates through dataset generation. ToxiCraft framework are illustrated in Figure 2.
@@ 3.1 Problem Statement
Robustness: In our setting, robustness is evaluated using cross-dataset evaluation, where data augmented on one dataset is used to evaluate performance on other datasets. Specifically, robustness is measured by the model’s performance consistency across different datasets $\mathcal{D}_{i}$, with data augmented from one dataset $\mathcal{D}_{j}$ and evaluated on another dataset $\mathcal{D}_{k}$. The variance of the model’s loss across these datasets captures the robustness:
$\text{Robustness}(\theta,\mathcal{D})=\text{Var}_{\mathcal{D}_{k}}(\ell(h_{\theta}(\mathcal{D}_{k}),y_{k}))$
where $h_{\theta}$ represents the model parameterized by $\theta$, $\ell(\cdot)$ is the loss function, and $\mathcal{D}_{k}$ are the instances from different datasets. Var captures the variance in performance across these datasets, reflecting the robustness of the model when trained on data from $\mathcal{D}_{j}$ and tested on others. We also provide a pseudocode in Appendix A.
@@ 3.2 ToxiCraft Framework
ToxiCraft Prompt Next, we added manually-set indicators to this data with Attribute Extraction by GPT (we call this process as ‘ToxiCraft Prompt’) and input 10% of the data extracted randomly from the seed data pool together with ‘ToxiCraft Prompt’ to GPT. Manually-set indicators include the intensification or weakening of tone, whether to increase swear words, whether to use irony, country and time. Each value will be randomly selected or masked. Thus, it is possible that none of the manually-set indicators are added, but at most only one answer will be selected from each indicator. For example, the country could be randomly set as the United States and the year as 2023.
@@
ToxiCraft effectively utilizes the Chain of Thought (COT) concept to gradually generate high-quality training data. Additionally, we compared the dataset generated by ToxiCraft with the dataset generated only using simple COT prompts, and detailed experimental results will be presented in Section 4.
@@ 4.1 Datasets and Processing
In this section, we outline the data employed in our experiments. We first introduce the four datasets containing harmful information in Section 4.1.1. Following this, we consolidate these datasets and present statistics in Section 4.1.2.
@@ 4.1.1 Harmful Information Datasets
It’s important to mention many datasets such as AIRAIT Waseem and Hovy (2016), only their tweet IDs and their labels were publicly available. Regrettably, a substantial portion of tweets (>= 40% positive datas) were inaccessible through the X API.
@@ 4.1.2 Data Processing and Splits
Our data preprocessing pipeline prioritizes identifying harmful content in English. Thus, we exclude any tweets in languages other than English and verify language using a Roberta-based language identifier. Additionally, to prevent overlap between similar datasets, we remove near-duplicated entries through normalization, ignoring duplicate entries and removing URLs and mentions. The majority of collected datasets focus on hate speech, and in some cases, offensive speech, which we consider harmful. For datasets with a binary hate classification task or a more detailed classification like CMS, where all "hate" subclasses are treated as one, we categorize them as harmful. Datasets focusing on specific types of hate speech, such as sexism, are also considered harmful. Moreover, datasets containing offensive speech are likewise classified as harmful in our case. Finally, in datasets like AHSD and HASOC, where a distinction between hate, offensive, or profound speech exists, we include entries labeled as hate speech or offensive, considering them as harmful. Entries labeled as normal or not-hateful are categorized as not harmful.
@@
Table 2: Comparison of Model Performances Across Different Databases and Data Generation Methods. The comparison parameter is MacroF1. Results outperformed training on gold label are underlined while the best performance on each dataset are bolded. Some baseline results come from Antypas and Camacho-Collados (2023).
@@ 4.3 Implementation Details
We employed GPT-4 OpenAI et al. (2024) as our agent model. For generations, in our experiment, we set N to be 1000, K to be 100, and M to be 3. For smaller-scale classification tasks, we utilized two language models of a reasonable size: BERT-base Devlin et al. (2019b) and RoBERTa-base Liu et al. (2019) as downstream models. The implementations provided by Hugging Face Wolf et al. (2020) were utilized for training and evaluating all language models. The fine-tuning parameters are listed in B, and we used macro-F1 to report the evaluation scores, ensuring consistency across models and tasks.
@@ 5 Experimental Results
Analysis on Robustness Figure 3 reveals ToxiCraft’s robust ability to generate synthetic data that enhances model generalization across datasets. In the AHSD to HASOC evaluation, we observe that as the seed size increases, the performance of models trained on ToxiCraft synthetic data rises significantly, peaking at a Macro F1 score of 67 for the 200 seed size. This surpasses the gold label performance for the AHSD dataset, which stands at a Macro F1 score of 54. Similarly, for the HASOC to HTPO evaluation, the upward trend continues with the performance of the 200 seed size reaching a Macro F1 score of 58, again outperforming the gold label performance, benchmarked at 56. These results not only demonstrate ToxiCraft’s capacity to produce high-quality synthetic data that captures the complexity of harmful content but also its effectiveness in adapting to different datasets. The ability to exceed gold label performance suggests that ToxiCraft could potentially reduce the need for extensive labeled datasets, particularly in specialized or sensitive areas like political content, where data acquisition is challenging.
6 Ablation Study
In this section, we present the results of our ablation study, which focuses on three key aspects: the choice of model, the impact of different modules, and the robustness of our framework under different settings. These experiments were conducted on the AHSD dataset, HTPO dataset, with 200 seed data and BERT as the downstream model.
6.1 Study on Backbone Model
In the early stages of our research, we experimented with several open-source models, including LLaMA2 13B, LLaMA3 8B and Flan-T5-XXL, as alternatives to GPT-4. However, the success rates and overall performance of these models were significantly lower than those achieved by GPT-4.
Table 4 summarizes the success rate and macro-F1 scores for each model. AHSD Dataset with 100 seed data is listed in C. GPT-4 clearly outperforms the alternatives, which justifies its selection as the preferred model for our experiments.
Model Success Rate (%) Macro-F1
GPT-4 78 89.0
LLaMA2 13B 32 38.7
Flan-T5-XXL 27 34.1
LLaMA3 8B 22 62.5
Table 4: Comparison of models on the AHSD dataset with 200 seed data and BERT as the downstream model.
6.2 Study on Modules
To assess the contributions of individual components in our framework, we conducted ablation studies by systematically removing key modules. The results, displayed in Table 5, highlight the significance of each module in model performance.
ToxiCraft Prompt: Removing the manually-set indicators in the ToxiCraft Prompt caused a notable decrease in macro-F1 score, from 89.0 to 85.2 on the AHSD dataset, and from 70.3 to 58.5 on the HTPO dataset, underscoring the importance of prompt tuning in the data generation process.
CAE: Excluding the CAE module reduced performance to 84.3 macro-F1 on the AHSD dataset and 63.8 on the HTPO dataset, demonstrating its crucial role in capturing contextual information.
TSR: Removing the TSR module resulted in the lowest performance (83.6 macro-F1 for AHSD, 65.1 for HTPO), further emphasizing its importance in improving model robustness and generalization.
Method AHSD HTPO
Full model 89.0 70.3
w/o ToxiCraft Prompt* 85.2 58.5
w/o CAE 84.3 63.8
w/o TSR 83.6 65.1
Table 5: Ablation study results on both the AHSD and HTPO datasets using MacroF1. * indicates ToxiCraft Prompt with manually set indicators.
6.3 Study of the Cross-Dataset Benefits of Modules
To further validate the robustness of our framework, we evaluated its performance on the HASOC dataset, using the AHSD dataset for data generation. As shown in Table 6, the full ToxiCraft framework exhibited the highest macro-F1 score (67.3), confirming its effectiveness in generating robust training data across datasets. When the manually set indicators in the ToxiCraft Prompt were removed, the macro-F1 score dropped significantly to 49.5, further underscoring the importance of these indicators in enhancing data diversity and performance. The exclusion of the CAE module resulted in a lower macro-F1 score of 50.3. Similarly, the absence of TSR led to a reduction in performance (57.4 macro-F1). In comparison, using only gold label data yielded a macro-F1 score of 54.7, showing the advantage of our method.
Method Macro-F1
Gold label 54.7
ToxiCraft 67.3
ToxiCraft Prompt w/ * 58.6
ToxiCraft Prompt w/o * 49.5
w/o CAE 50.3
w/o TSR 57.4
Table 6: Robustness study on the HASOC dataset using data generated from the AHSD dataset, * means manually set indicators.
7 Managing Risks of Generating Hateful Content
The use of large language models (LLMs) to generate hateful content presents significant ethical concerns, particularly around amplifying harmful biases. To mitigate these risks, we implemented strict content review processes to filter inappropriate outputs and limited the scope of generation to controlled experimental environments. We carefully ensured that synthetic data remained focused on research objectives. All experiments adhered to ethical guidelines, underwent reviews, and complied with responsible AI standards. The datasets generated were made available with clear warnings to promote transparency and encourage responsible use in future research. We will actively collaborate with the community to monitor, responding swiftly to any indications of misuse.
@@ 9 Future Work
Future research for ToxiCraft will focus on three main objectives: 1) enhancing multilingual capabilities, 2) refining the seed data selection process, and 3) reducing reliance on high-cost LLMs. For multilingual data generation, future efforts will explore translating harmful content from English into other languages while preserving both local nuances and the target language characteristics. In terms of optimizing seed data selection, the focus will be on refining the methodology to identify the most diverse and representative examples within a dataset. One proposed approach is to analyze word embeddings and select seed data points whose vectors are the least similar, ensuring broad coverage of the data space. Lastly, cost-effective model exploration will involve identifying more affordable and transparent alternatives to GPT-4, such as the Mistral model, for generating high-quality synthetic samples of harmful content Hui et al. (2025).
@@ Appendix A Pseudocode for ToxiCraft
Appendix B Training parameter
We fine-tuned the learning rate, warm-up steps, number of epochs, batch size, and other hyperparameters for each model. For BERT, we use learning rate=$1.8282\times 10^{-5}$, epoch=3, batch=4, warm-up step=30.And for RoBERTa, we use learning rate=$1.1856\times 10^{-5}$, epoch=3, batch=4, warm-up step=30. All fine-tuning experiments were conducted using A40 and GTX 4080 Super GPU, and the results are reported in terms of macro-F1 score.
Appendix C AHSD Dataset with 100 seed data Success rate
Table 7 presents the performance of different models on the AHSD dataset, using 100 seed data and evaluated with BERT as the downstream model.
Model Success Rate (%) Macro-F1
GPT-4 76 74.7
LLaMA2 13B 42 10.8
Flan-T5-XXL 33 16.1
Table 7: Comparison of models on the AHSD dataset with 100 seed data and BERT as the downstream model.
Appendix D Example of ToxiCraft framework
In Figure 4, we provide a example of our ToxiCraft to help better understand.
Figure 4: Our ToxiCraft Framework example